# Zhiyuan Liu3, Jinan Xu1, Yuhua Li2, Ruixuan Li2, Maosong Sun3

### v2

@@ 7.3 Effects of Different Data on Multi-Image Understanding
As observed in Table 6, Migician shows an improvement in multi-image understanding. We further conduct an ablation study to analyze the effects of different data subsets. Specifically, we train two models with either the multi-image grounding and multi-image understanding data removed from the training set.
The results in Table 6 reveal that grounding data generally aids multi-image understanding. In 4 out of 5 benchmarks, the full dataset achieves the highest performance compared to models trained with any subset of data removed. In contrast, directly fine-tuning with only general data does not consistently lead to a performance boost. However, when combined with fine-grained grounding data, the model experiences a notable improvement.
@@ 8 Conclusion
Despite our comprehensive discussion of the MIG challenge, there still remain several limitations. First, due to the computational budget, we haven’t verified the effectiveness of our training methods on larger 70B scale models. Secondly, in spite of intensive grounding training, our model is still confronted with inaccurate grounding issue, especially in complicated or messy scenarios. Lastly, our training methods and benchmark construction mainly focus on the REC task. Although Migician possesses decent REG capacity, this topic is still insufficiently discussed.
@@ A.1 Spontaneous Grounding
Our benchmark evaluates the spontaneous grounding through three distinct tasks below, which aim at assessing model’s ability to autonomously discover insidious connections across various images and accurately recognize then locate the target.
@@
Given two similar images with a single subtle difference, the model is instructed to recognize and ground this difference in the second image, requiring keen perceptual skills.
@@
It refers to automatically recognizing and grounding the common object appearing in all images within an image group, which in our bench, each shares one definite common object.
@@
Models must focus on the primary difference between two images captured from slightly different perspectives, ignoring other minor variations caused by shifts in the viewpoint.
@@ A.2 Reference Grounding
This challenge, which mainly includes Group Grounding, tests a model’s ability to link a textual reference to a target object within its corresponding specific image. Given a set of images and one textual query, the model must identify the correct image then accurately ground the target object within it.
@@
These tasks examine model’s ability to effectively utilize visual reference information and incorporate it into the searching process.
(1) Visual Referring Grounding. In this task, a pair of images is provided—a source image with a clear object and a target image containing multiple elements. The model must locate the referenced object in the target image.
@@
(3) Object Tracking. This task involves tracking a target object across a sequence of video frames. The object is highlighted with a red bounding box in the first frame, and the model must follow it throughout the sequence.
(4) Multi-view Grounding. Here, the model must locate the same target object across multiple images taken from distinct viewpoints.
@@
These tasks combine information from both modalities to assess cross-modal reasoning abilities.
(1) Correspondence. The model must ground semantically or functionally similar regions within a target image. This finer-grained task focuses on object regions rather than whole objects, demanding an in-depth understanding of visual semantics.
@@
Our comprehensive benchmark offers a rich, multi-faceted evaluation across various real-world scenarios and domains, extending beyond simple image pairs to include longer and more complex image contexts. By ensuring that each task is well-defined and unambiguous, we facilitate objective and definitive assessments..
Appendix B Single-Image CoT Failure Patterns
As shown in Figure 5, the four representative failure patterns are (a) special multi-image format, (b) abstract visual information, (c) CoT error propagation, (d) step-2 inference error.
When the multiple images are formatted in a special pattern, where our target object is missing in the target image, like, the information in this image is insufficient to perform grounding.
The abstract visual information refers to situations where the intricate visual cannot be adequately converted in textual description to perform accurate grounding. In Figure 5, the simple description "a close-up of a woman’s face" cannot distinguish which face the target is in Image-1.
Each reasoning step of CoT could be incorrect, which could potentially leads to the error propagation issue [44]. In Figure 5, the conclusion draw from the first step is incorrect, which directly leads to the mistake in the second step.
The last failure pattern refers to cases where the erroneous reasoning step appears at the second step, failing to accurately ground or follow instruction.
Appendix C MGrounding-630k Data Curation Details
C.1 Transforming Existing Data
@@
Describing the differences of the two nearly same pictures is a well discussed topic, yet they focus on the coarse-grained semantic feature, failing to precisely recognize the part of differences. After comprehensive survey on this area, we have collected high-quality and fully labeled image difference data from Spot-the-diff [16], Img-diff [18], MagicBrush [50] and CLEVR-change [32].
During the construction process, we ensure the diversity of the content by (1) incorporating numerous prompt formats generated by GPT-4, (2) constructing CoT process to assist the model gradually and progressively reaching the final answer, while also fully utilizing the annotation available in the dataset.
@@
Recognizing and grounding the main common object in multiple images is an interesting yet non-trivial task for models, which firstly requires them to simultaneous look at multiple images, disentangle the common object, then finally grounding the target object in every single image. We take diverse data sources from ImageNet [6], COCO [26] and Objects365, where the annotations are abundant and rich. Through their annotations, we group the images that share the same object together with a threshold of the proportion it takes for the whole image to filter out too tiny objects. We empirically find such threshold effective at eliminating ambiguity where there could be multiple common object candidates, which results in clear and definite training examples. We further reduce ambiguity by skipping inappropriate classes where there are always multiple possible candidate such as couch, dinning table, keyboard and etc.
@@
The original object tracking emphasize more on tracking the target object in a video sequence, which resembles with multi-image sequences. As a well discussed topic, we select the large scale TrackingNet [31], LaSOT [8], GOT-10K [13] and MOT-2017 [30] datasets as our data sources. During the dataset construction process, we have simplified the original long sequence to 4-6 image per training example, maintaining its original core feature while keeping efficiency. For each image, we extract the key frame with an appropriate interval to ensure the obvious movement of the key target. We also involve small proportion of the ordering judge for continuous images to enhance the model’s temporal understanding ability.
@@
This part of training data is designed to imitate finding the object of the source image in the target image. We mainly utilize the ImageNet-2012 subset to construct image pairs, with the first one taking large proportion of the whole image, and the second one containing smaller target object that may take efforts to spot. In total, the referring grounding dataset covers a wide range of objects that is beneficial for the model to generalize.
@@
Conventional visual grounding is mostly limited to single image context, while in real world scenario, we often need to recognize the target object from a messy piles of pictures. Group Grounding, which locates the target among a group of different images, is the exact task to fill in such gap and enrich the versatility of traditional grounding. When constructing this part of data, we take advantage of the single image GranD rec and reg conversation data [34] with the quantity of 3M. With further filtering and combining 3-5 images per group, we finally obtained a collection of 12w high-quality training data for stage-1 training(grounding injection training), which is effectively at enhancing the co-reference and image-level locating ability of models.
@@
Region locating refers splitting several pieces of semantic-rich region from the source image and then recognize the exact location of this region image in the source picture. To guarantee extracting meaningful regions, we utilize the Objects365 dataset and extract the bounding box area as the regions. To further improve quality, we set a series of filtering mechanism:(1) content richness: we select the images that have more than 10 bounding box annotations to avoid too plain or simple cases. (2) aspect ratio: we keep the aspect ratio between 0.5-2 to avoid the excessively thin bounding box areas that models may fail to effectively tackle. (3) size: we keep the region ratio of the whole image between 0.2-0.49 and an absolute pixels above 2000 to avoid excessively tiny and obscure region images. Noticeably, due to our meticulously crafted mechanism and the feature of this task, the resulting training data encompasses the cases of person recognition in the source image, analogous objects distinguishing and tiny details recognition that are non-trivial even for human.
@@
C.2 Synthesizing Free-form MIG Data
The algorithm for CLIP adaptive similarity image input is shown in Algorithm 1. We further display our prompt template for image caption generation, bounding box label refinement and instruction tuning data generation in the following pages.
@@ C.2 Synthesizing Free-form MIG Data
Appendix D Details of Two-Stage Training
@@
Appendix E Evaluation Implementation
When directly requiring the model to generate bounding box coordinates for each image, due to their limited multi-image grounding ability and insufficient instruction following ability, the answer obtained in this way is largely unfaithful and mostly unsatisfactory in instruction following, failing to objectively reflecting the real grounding ability of the model.
@@ Appendix E Evaluation Implementation
The performance of Migician is presented in Table 5. Although mainly targeted at multi-image grounding, Migician still maintains well on conventional single-image grounding task.
@@
V*-Bench Implementation
When adapting a single high-resolution image to a multi-image setting, we require the model to first ground, then judge the attributes. We start by dividing the original image into four equally sized sub-images. This transforms the single-image localization problem into a group grounding task, where the model is required to perform both image-level and object-level grounding. The model generates outputs such as: "The tiny dog is located in Image-4, (346,763),(431,835)."
Once the object is localized, we send only the identified sub-image and its corresponding region of interest to the model for attribute recognition and relative relationship analysis. To ensure contextual completeness, we expand the extracted region by slightly enlarging its width and height, adding an additional 50 pixels to each dimension.
Appendix F Multi-Task Learning
@@
Appendix G Case Study

### v3

@@ 7.2 Effects of Different Data on Multi-Image Understanding
As observed in Table 3, Migician demonstrates an improvement in the multi-image understanding. We further conduct an ablation study to analyze the effects of different data subsets. Specifically, we train two models with either MGrounding-630k or multi-image understanding data removed from the original training set. The results in Table 6 reveal that grounding data generally contributes to multi-image understanding. In 4 out of 5 benchmarks, the full dataset achieves the highest performance compared to models trained with any subset of data removed. In contrast, directly fine-tuning with only general data does not consistently lead to a performance boost. However, when combined with fine-grained grounding data, the model experiences a notable improvement.
7.3 Performance across Difficulty Levels
To comprehensively assess Migician across varying scenario complexities in MIG-Bench, we have established a three-tier difficulty classification (Easy/Medium/Hard) through joint consideration of three key factors: (1) the number of input images, (2) the accuracy rates of four representative baselines (Qwen2-VL, InternVL2, and their CoT variants), and (3) the average IoU improvement when applying CoT. Specifically, an instance is classified as Easy if either (a) more than two models achieve correct answers with fewer than four input images, or (b) the CoT-enhanced models demonstrate an IoU improvement exceeding 0.15. Conversely, instances are deemed Hard when no more than one model succeeds despite processing over four input images. All remaining cases that fall between these thresholds are categorized as Medium difficulty.
Through this way, we can evaluate model capabilities across diverse scenario complexities. As shown in Table 7, Migician demonstrates remarkable performance advantages across all difficulty levels while exhibiting particularly strong capabilities in challenging scenarios. Noticeably, the performance gap widens significantly in medium and hard difficulty settings as Migician achieves approximately 30% greater accuracy compared to the baseline models. This pronounced advantage in complex scenarios highlights Migician’s superior capacity for free-form MIG.
@@ 8 Conclusion
Despite our comprehensive discussion of the MIG challenge, several limitations remain. First, due to computational constraints, we have not verified the effectiveness of our training methods on larger 70B-scale models. Second, the current model can still produce hallucinated outputs sometimes as conventional MLLMs. Lastly, our training methods and benchmark construction mainly focus on the REC task. Although Migician possesses decent REG capacity, this topic is still insufficiently discussed.
@@ Appendix A Benchmark Tasks Definition
Dataset Images Ave-I Max-I Multi-Image Multi-Task Instance-Labeled Instances Reference
Q-Bench 3489 2.0 2.0
Mantis-Eval 542 2.5 5.0
BLINK 3612 1.9 4.0
MIRB 3497 3.8 42.0
Refcoco/g/+ 3900 1.0 1.0 7596 T
HC-Refcoco/+/g 1521 1.0 1.0 3754 T
GigaGround 3775 1.0 1.0 61353 T
MC-Bench 3345 2.0 2.0 3202 T
MIG-Bench 5887 3.8 6.0 4295 T/I/T+I
Table 8: Comparison of MIG-Bench with other benchmarks.
@@ A.1 Spontaneous Grounding
Our benchmark evaluates spontaneous grounding through three distinct tasks below, which aim at assessing model’s ability to autonomously discover insidious connections across various images and then accurately locate the target.
@@
Similar to the spot-the-difference puzzle, given two similar images with a single difference, the model is instructed to recognize and ground this difference in the second image, requiring simultaneous and keen perception of both images.
@@
It refers to automatically recognizing and grounding the common object appearing in all images within an image group, which shares one definite common object in our benchmark.
@@
Models must focus on the primary difference between two images captured from slightly different perspectives, ignoring other minor variations caused by shifts in the viewpoint. The incorporation of view changes presents a greater challenge for the model and better reflects real-world scenarios, where variations in perspective are inevitable.
@@ A.2 Reference Grounding
This challenge, which mainly includes Group Grounding, tests a model’s ability to link a textual reference to the target object within its corresponding image among an image group. Given a set of images and one textual query, the model must firstly identify the correct image then accurately ground the target object within it, additionally incorporating image-level locating compared with conventional grounding task.
@@
These tasks focuses on effectively utilizing visual reference information and incorporate it into the locating process.
(1) Visual Referring Grounding. In this task, a pair of images is provided—a source image with a clear object and a target image containing multiple elements. The model must perceive the referenced object of the source image and then locate it in the target image.
@@
(3) Object Tracking. This task involves tracking a target object across a sequence of video frames. The object is highlighted with a red bounding box in the first image, and the model must keep track of it throughout the entire sequence.
(4) Multi-view Grounding. Here, the model must locate the same target across multiple images taken from distinctive viewpoints, for instance egocentric view and exocentric view.
@@
These tasks combine information from both modalities to assess cross-modal reasoning&grounding abilities.
(1) Correspondence. The model must ground semantically or functionally similar regions within the target image. This finer-grained task focuses on object regions rather than whole objects, demanding an in-depth understanding of visual semantics.
@@
Our MIG-Bench offers a rich, multi-faceted evaluation across various real-world scenarios and domains, extending beyond simple image pairs to include longer and more complex image contexts. By ensuring that each task is well-defined and unambiguous, we facilitate objective and definitive assessments.
Appendix B Comparison of MIG-Bench with Other Benchmarks
A detailed comparison with other benchmarks are provided in Table 8. Current evaluations for MLLMs primarily focus on single-image perception, understanding, reasoning, or grounding (e.g., MME, MMBench, Refcoco), leaving the multi-image scenario largely unexplored. While benchmarks such as Mantis-eval, BLINK, and MIRB are representative of multi-image evaluation, they concentrate on image-level and general understanding of multiple images, failing to comprehensively assess a model’s fine-grained grounding skills in the multi-image scenario.
Unlike traditional grounding benchmarks or multi-image benchmarks, MIG-Bench introduces the grounding task into a multi-image scenario, covering a series of 10 distinct tasks. The most relevant benchmark to our work is MC-Bench, a contemporaneous study. MC-Bench evaluates the multi-context grounding capabilities of MLLMs by asking them to accurately locate the corresponding object based on a text prompt in the correct image from a given image pair. However, it has limitations in terms of the fixed number of input images and the restricted forms of queries. In contrast, the proposed MIG-Bench offers more flexible task formats, focusing on evaluating models’ abilities in free-form multi-image understanding.
Appendix C Single-Image CoT Failure Patterns
We have analyzed more failure patterns of the CoT framework in Figure 5, categorized into perceptual and reasoning flaws.
For the former, the framework falls short when multiple images are organized in a manner where only integrating all their visual information could tackle MIG (i.e. finding the location of missing people in the second image), or when the textual content could not sufficiently represent the visual information.
Regarding reasoning errors, inaccuracies can arise at various stages of the reasoning process, undermining the framework’s overall accuracy and effectiveness.
These failure patterns highlight the significant limitations of simply integrating the different capabilities of current models through a simple CoT framework, underscoring the need for an end-to-end model capable of directly performing the MIG task.
Appendix D MGrounding-630k Data Curation Details
D.1 Transforming Existing Data
@@
Describing the differences among two nearly identical pictures is a well discussed topic, yet previous attempts capture the differences through textual phrases, failing to precisely recognize their locations. After a comprehensive survey on this area, we have collected high-quality and fully labeled image difference data from various existing datasets: Spot-the-diff [17], Img-diff [19], MagicBrush [53] and CLEVR-change [34].
By these collected datasets inherently contain much noise in them, for instance, inaccurate difference caption, incorrect bounding box coordinates and etc. We filter the inaccurate bbox labels from Spot-the-Diff and only preserve the correct ones and their difference captions. Additionally, we conduct down-sampling on Img-Diff due to its diffusion generation [37] based nature and consequent inaccuracies.
During the construction process, we ensure the diversity of the content by (1) incorporating numerous prompt formats generated by GPT-4 and improving the instruction-following ability of the model, (2) constructing CoT process to assist the model gradually and progressively reaching the final answer by fully utilizing the annotation available in the dataset.
@@
Grounding the primary common object across multiple images is a challenging task for models. It requires them to simultaneously perceive multiple images, isolate the common object, and then accurately ground the target within each image.
In our approach, we leverage diverse data sources, including ImageNet [6], COCO [27], and Object365, which are rich in extensive annotations. To organize the images, we group those containing the same object based on object labels, applying a threshold to filter out objects that occupy too small a proportion of the image. This threshold, determined empirically, effectively mitigates ambiguity when multiple candidate objects could be considered the common object, thereby producing clear and definitive training examples. We further reduce ambiguity by removing classes that often co-occur with other object, for instance keyboard, knife, couch, dinning table and etc.
@@
The multi-image setting is well-suited for the object tracking task, which involves both temporal and spatial grounding. To build our dataset, we have carefully selected large-scale, well-annotated datasets including TrackingNet [33], LaSOT [8], GOT-10K [14], and MOT-2017 [32]. During dataset construction, we extract 4-6 images from each original sequence at specific intervals, preserving key features while maintaining efficiency. Additionally, we introduce a small proportion of temporal order judgment data from continuous video frames to strengthen the model’s temporal reasoning capabilities.
@@
This training data simulates the process of locating an object from a source image in a target image. We primarily use the ImageNet-2012 dataset to create image pairs, where the source object is fully visible in the first image, and the target object is hidden in the second. Empirically, this design generates challenging training examples that require the model to first recognize the source object and then search for it in the target image.
@@
Conventional visual grounding is typically limited to single-image contexts. However, in real-world scenarios, it is often necessary to recognize a target object within a cluttered collection of images. Group Grounding addresses this limitation by enabling the identification of the target among a group of images, thereby enhancing the versatility of traditional grounding methods.
For the construction of Group Grounding training data, we leverage the large-scale GranD rec and reg conversation dataset [36], which contains 3 million examples. After filtering out noisy data and grouping 3-5 images per set, we curate a high-quality collection of 120k instances for stage-1 training. This dataset effectively enhances the model’s ability to perform image co-reference [18], image-level grounding, and instance-level localization.
@@
Region locating involves slicing an image into several semantically rich regions and identifying the precise locations of these regions within the source image. To extract meaningful regions, we use the Objects365 dataset, selecting labeled bounding box areas as the regions of interest. To further enhance the quality of the regions, we apply a series of filtering criteria: (1) Content Richness: We select images with more than 10 bounding box annotations to avoid overly simple or plain cases. (2) Aspect Ratio: We retain regions with an aspect ratio between 0.5 and 2 to exclude excessively narrow bounding boxes that may be difficult for the model to handle. (3) Size: We ensure that the region-to-image ratio lies between 0.2 and 0.49, with an absolute pixel count above 2,000, to exclude tiny and obscure regions that may lack sufficient detail. Notably, due to our carefully designed filtering mechanism and the inherent characteristics of the task, the resulting training data predominantly includes cases of person recognition, distinguishing between similar objects (i.e. chairs, bowls, cars and etc), and recognizing tiny details—tasks that are non-trivial even for humans.
@@
D.2 Synthesizing Free-form MIG Data
The algorithm for CLIP adaptive similarity image input is shown in Algorithm 1. We further display our prompt template for image caption generation, bounding box label refinement and instruction tuning data generation in the following pages and several stage-2 data examples in Figure 8.
@@ D.2 Synthesizing Free-form MIG Data
Appendix E Details of Two-Stage Training
@@
Appendix F Evaluation Implementation
Polling-based Evaluation
When directly requiring the model to generate bounding box coordinates for each image, due to their limited multi-image grounding ability and insufficient instruction following ability, the answer obtained in this way is largely unfaithful and mostly unsatisfactory in instruction following, failing to objectively reflecting the real grounding ability of the model. . Empirically, instructing the models to directly generate all bounding box coordinates results in very serious instruction following issue. The models struggle with plausible outputs, with their answers mostly containing only one bounding box or pure text analysis.
@@ Appendix F Evaluation Implementation
V*Bench Evaluation Implementation
We transform the high-resolution single image grounding task into a MIG challenge. Specifically, we slice a single high-resolution image into multiple sub-images and directly convert the problem into the group grounding task, which first requires the model to perform image-level locating and then ground the target in that specific image. By utilizing the MIG ability of Migician, we can locate the regions relevant to the input question. Afterward, the model combines the identified region with the original image to generate the answer for the input question, achieving high-accuracy results.
@@
Single-Image Grounding
Human-Level Evaluation
We engaged five human volunteers to answer questions from MIG-Bench. For the static image difference, common object grounding, region locating, and group grounding tasks, we randomly selected 20% of the test examples for efficient evaluation.
The volunteers are instructed to answer the MIG questions by directly drawing bounding boxes on the images, with the resulting annotations then compared to the ground truth using automatic IoU calculation. To ensure unbiased results, evaluations from different volunteers were conducted separately, preventing mutual distraction or information leakage.
Appendix G Multi-Task Learning
@@
Appendix H Case Study